# Probabilistic operator learning: generative modeling and uncertainty quantification for foundation models of differential equations

### v1

@@ Abstract
Key words. Probabilistic operator learning; Generative modeling; Uncertainty quantification; In-context learning, Bayesian inference; Generative ICON
@@ 1 Introduction
Many operator learning methods, including FNOs and DeepONets, are trained using supervised learning, where a mapping is learned from given input data that contains model parameters along with their initial and boundary value data to given output data that consist of solutions [14]. In contrast, ICON is trained to learn context, where the model is not explicitly trained on parameters but instead rely only on example initial-boundary value data and corresponding solution pairs. The ICON architecture is based on transformers, the neural network architecture that underpins large languages models [15, 16], which have shown to be capable of solving multiple types of tasks given a particular context defined by example ‘condition-solution’ pairs. ICON is an example of a foundation model for differential equations as it is pre-trained on a variety of ODE and PDE problems, and then fine-tuned to learn specific solution operators given only example pairs of initial-boundary value data and solutions. Our probabilistic formulation of ICON provides a model for understanding the mathematics of fine-tuning of foundation models for differential equations. This ability to fine-tune ICON given new example pairs to learn solution operators of unseen differential equations is why [3, 4] describes it as an operator learner rather than an operator approximator: it adapts on-the-fly to new contexts, learning the correct operator without being given explicit information about the underlying model parameters.
@@
To understand ICON, operator learning, and foundation models of differential equations more generally, we present the notion of probabilistic operator learning, which requires modeling the data generating process that produces the training datasets. For operator learning, we model this data generating process through the formalism of random differential equations (RDEs) [17], upon which we may understand the connections of operator learning and ICON to conditional expectations, least-squares estimators, and Bayesian inference. Random differential equation provide a natural mathematical framework for understanding not just ICON, but many foundation models trained on data from many different differential equations and initial and boundary data. The probabilistic setting also provides a rigorous way of describing confidence in model predictions by characterizing model uncertainty.
@@
ICON learns a function that maps example pairs of input-output data of a differential equation along with a new unseen input, and maps it to the predicted output. This is in contrast to traditional operator learning methods in which a function that maps inputs and operator parameters to solution outputs. Our primary contribution is showing that ICON directly approximates the expected value of a posterior predictive distribution, bypassing an explicit characterization of a Bayesian posterior distribution, meaning that ICON performs Bayesian inference implicitly. In other words, ICON infers the appropriate operator parameters given the example pairs and then applies the learned operator to the new inputs to predict the corresponding output in a single step. In-context learning more broadly has been understood as approximating the posterior predictive distribution [18], although this interpretation has only been studied in the context of large language models.
@@
With the knowledge that ICON approximates the mean of the posterior predictive distribution, our second main contribution is adapting ICON to generate from the posterior predictive distribution. In the context of ICON, the posterior predictive characterizes the distribution of solutions to the differential equation conditioned on example data, meaning that its statistics describes the uncertainty of the predictions as well. Tools from conditional generative modeling are needed to generate from the posterior predictive, and we show that being able to produce samples from it provides a way to quantify uncertainty in ICON’s predictions. Generating from the posterior predictive distribution is an example of conditional generative modeling. When training ICON, or foundation models for differential equations more broadly, we are given input-output pairs of conditions and solutions which comes from some joint probability measure. Conditional generative modeling takes input-output pairs to characterize the distribution of outputs given unseen inputs. ICON takes this a step further where it can characterize the distribution of solutions given new conditions and input-output data that inform the model parameters.
@@ 1.1 Contributions
  • Using the RDE formalism, we explain in-context operator learning as implicitly performing Bayesian inference, in which ICON approximates the mean of the posterior predictive distribution. In Section 3, we describe in-context learning in probability spaces with Banach and Hilbert space valued random variables. Since solutions of differential equations reside in function spaces, it is natural — and indeed, necessary — to adopt this setting as model prediction will also reside in these spaces. This functional analytic framework provides the foundation for understand when least-squares predictors coincide with conditional expectations in operator learning–based approaches to differential equations.
  • We present a generative form of in-context operator learning that is able to generate from the posterior predictive distribution. This is a form of uncertainty quantification, which describes the level of confidence ICON has in prediction. Building on the abstract formulation in Section 3, we establish the existence and properties of a generative ICON method in Section 4, as well as its ability for performing uncertainty quantification. Numerical examples that demonstrate our approach is discussed in Section 5.
@@
As we discussed in Section 1, probabilistic operator learning requires us to define the data generating process that produces the training data for operator learning. Random differential equation provides the probabilistic model that generates this training data, and provides the proper formalism for which many operator learning methods can be understood. In particular, we highlight the probability measures that are induced by random ODEs, boundary value problems, and time-dependent PDEs. Solutions to differential equations are elements of function spaces, i.e., Banach and Hilbert spaces, and therefore, the probability measures we study are defined on Banach and Hilbert spaces.
@@ 2.2 Random ODEs
where $u(t,\omega)\in\mathbb{R}$ is a function of time, $c(t,\omega)$ is a Gaussian process, and $\gamma_{1}(\omega)$, $\gamma_{2}(\omega)$, and $u_{0}(\omega)$ are $\mathbb{R}$–valued uniform random variables. The random function $a(t,u,\omega)=\gamma_{1}(\omega)c(t,\omega)+\gamma_{2}(\omega)$ is parametrized by $\gamma_{1},\gamma_{2}$, and $c(t,\omega)$. In practice, Gaussian processes are approximated by finite dimensional Gaussian vectors $\mathbf{c}\in\mathbb{R}^{\mathbf{N_{c}}}$, and so the the random function $a$ is modeled by a finite dimensional random vector. The probability measure that arises from (4) can be represented in terms of a density $Q^{N}$
@@ 2.3 Random boundary value problems
Here we consider random boundary value problems (BVPs). The solution no longer evolves in time, however $V$ and $W$ are function spaces defined over some domain $E\subset\mathbb{R}^{d}$, for example, $V=L^{2}(E)$ and $W=L^{2}(\partial E)$ Generally, we may consider equations of the form
@@
where $\mathbf{u}(\omega)\in\mathbb{R}^{N}$, $\mathbf{u}_{i}(\omega)=u(x_{i},\omega)$, and $\text{d}\mathbf{W}=\,\text{d}a\,\text{d}\mathbf{c}\,\text{d}\mathbf{k}\,\text{d}u_{0}\,\text{d}u_{L}\,\text{d}\mathbf{u}$. Solutions to boundary value problems are typically approximated by finite difference or element schemes, where the solution at a particular stencil point or element is explicit dependent only on a local neighborhood of the point or element. In this example, one will likely use a central difference scheme so that the density can be expressed with the following conditional independence structure
@@ 3.2 Probability measures of in-context operator networks
To clarify the distinction between ”classical” operator learning methods — such as DeepONets [2], Fourier neural operators [1], and neural operators more broadly [23]— and in-context operator learning, we briefly discuss the former in our probabilistic context. Classical operator learning is also uses training data of the form (13) with $J=1$, but the key difference is that the operator parameters $\alpha_{m}$ are observed during training and serve as inputs to the operator. Formally, classical operator learning seeks a mapping
@@ 3.3 In-context operator learning implicitly performs Bayesian inference
ICON is an example of amortized and likelihood-free inference in a distinct way: rather than repeatedly recomputing or even explicitly characterizing the posterior distribution, it maps observations directly to the posterior predictive distribution, bypassing the bottleneck of computing the posterior entirely. This is an advantage when one simply wants a predictive model without the need to fully characterize the posterior. Posterior distributions are often difficult to characterize particularly when they are supported on a low-dimensional manifold, are highly convex, which makes them difficult to sample. In Section 5, we consider an ODE that exhibits non-identifiability, meaning that multiple parameters may produce the same observational data. The resulting posterior then lies on a lower-dimensional manifold that is nontrivial to characterize. ICON bypasses all of these challenges by approximating the mean of the posterior predictive directly. We summarize the Bayesian explanation of deterministic in-context operator learning [3, 4] in Figure 2. In Section 4, a generative formulation of ICON will be able to characterize uncertainties of the predictive model.
@@ 4.1 Posterior predictive sampling via generative ICON
We emphasize that Theorems 4.1 and 4.2 is agnostic to the specific choice of generative modeling algorithm. While we train GenICONs using GANs in our numerical examples in Section 5, one may consider other generative models on function spaces [37, 38, 36, 35]. In Section 4.2 we provide algorithmic details describing the training of a conditional GANs from data generated for ICON.
@@
Generative ICON provides uncertainty quantification without additional cost by characterizing the distribution of predictions. By repeatedly sampling from the reference distribution with the same condition $y^{J}$ and context, GenICON produces an empirical distribution for the posterior predictive distribution from which we can estimate the predictive mean, higher moments, as well as distributions of other observables. Like ICON, GenICON is a likelihood-free and amortized inference method that can produce posterior predictive samples given context without the need to recompute the posterior distribution each time there are new observations. Figure 3 illustrates the capabilities of GenICON in comparison to standard ICON. ICON only produces the posterior predictive mean, which is a single function, while GenICON produces samples which can be used to not only characterize the mean but also the uncertainty around the predictive mean. In a sense, GenICON can be viewed as a generalization of Gaussian processes, which are used to make new predictions given only a new input and data given in ordered pairs [39].
Figure 3: Illustration of the capabilities of ICON vs generative ICON. Standard ICON returns a single solution, which represents the approximation of the mean of the Bayesian posterior predictive distribution. Generative ICON characterizes uncertainty of the mean solution through samples from the posterior distribution, which is represented by the gray paths. The blue region represents the the variance of the posterior predictive around the mean solution.
@@
If we are only interested in the first moment of (19), i.e. the conditional expectation (20), then we do not need to consider a full probabilistic formulation, as Theorem 3.1 shows, and in this case the least squares regression (16) sufficient. If however, we need probabilistic information, such as uncertainty quantification bounds for predictive guarantees for the conditional expectation predictions, then we need the full probabilistic formulation. The full probabilistic formulation is necessary for numerous other critical tasks, e.g. formulating and solving inverse problems using ICON, or carrying out model fine-tuning with UQ capabilities.
@@ 4.2 A generative ICON model based on generative adversarial networks
We detail one approach for constructing a generative model for sampling from the posterior predictive distribution (25). Suppose we are given the training data in the form of (16). To simplify presentation, denote the context, i.e., example condition-QoI pairs, by $\mathbf{c}=\{y^{j},z^{j}\}_{j=1}^{J-1}$. Let $\mathcal{G}_{\theta}:H\times(Y\times Z)^{J-1}\to Z$ be a family of conditional generators parameterized by $\theta$ that maps reference samples $\eta\sim\mathbb{P}_{\eta}$ and context $\mathbf{c}$ to an output sample $z^{J}_{\theta}=\mathcal{G}_{\theta}(\eta,\mathbf{c})$. The map induces a conditional measure $\mathbb{Q}_{z^{J}_{\theta}|\mathbf{c};\theta}(\cdot|\mathbf{c})=\mathcal{G}_{\theta}(\cdot,\mathbf{c})_{\sharp}\mathbb{P}_{\eta}$, as well as a parametrized family of joint measures $\mathbb{Q}_{\mathbf{c},z^{J}_{\theta};\theta}(\cdot)=\mathbb{Q}_{z^{J}_{\theta}|\mathbf{c};\theta}(\cdot|\mathbf{c})\otimes\mathbb{Q}_{\mathbf{c}}(\cdot)$, which will approximate the joint measure $\mathbb{Q}^{J}(\cdot)$ in (18).
@@ 5 Numerical examples
We demonstrate the Bayesian interpretation of ICON as well as our novel generative ICON through some simple numerical examples. Sections 5.1 and 5.2 demonstrates that ICON is robust to degenerate posteriors, model non-identifiability, and partial and noisy observations. These properties can be easily explained through a Bayesian inference. Sections 5.3 and 5.4 demonstrate generative ICON and its ability to capture the posterior predictive distribution. Examples that use ICON use the same model architecture as described in [3], while the generative ICON model architecture and training is described in Appendix A.
@@ 5.1 Robustness of ICON to degenerate posteriors and non-identifiability
We consider variations of the forward and inverse problem for Example 1 through the Bayesian interpretation of ICON. In consider the same random ODE under two different parametrizations:
@@
We let the parameters be the coefficients: $\alpha(\omega)=(\gamma_{1}(\omega),\gamma_{2}(\omega))\in\mathbb{R}^{2}$ and $\alpha(\omega)=(\eta_{1}(\omega),\eta_{2}(\omega),\eta_{3}(\omega))\in\mathbb{R}^{3}$. In the forward problem, the conditions are $y(\omega)=(c(t,\omega),u_{0}(\omega)\in L^{2}(0,T)\times\mathbb{R}$ and the quantity of interest be $z(\omega)=u(t,\omega)\in L^{2}(0,T)$. In the inverse problem, the roles are reversed — $y(\omega)=u(t,\omega)\in L^{2}(0,T)$ and $z(t,\omega)=c(t,\omega)\in L^{2}(0,T)$. For all cases, we let $T=1$. Notice that in the 2-parameter problem, the parameter are exactly identifiable with only a single demo (example-QoI pair). Given $(c(t,\omega),u(t,\omega)$, the parameters $(\gamma_{1},\gamma_{2})$ are determined by solving the linear system

### v2

@@ Abstract
Key words. Probabilistic operator learning, Generative modeling, Uncertainty quantification, In-context learning, Bayesian inference, Generative ICON
@@ 1 Introduction
Many operator learning methods, including FNOs and DeepONets, are trained using supervised learning, where a mapping is learned from given input data that contains model parameters along with their initial and boundary value data to given output data that consist of solutions [14]. In contrast, ICON is trained to learn context, where the model is not explicitly trained on parameters but instead relies only on example initial-boundary value data and corresponding solution pairs. The ICON architecture is based on transformers, the neural network architecture that underpins large languages models [15, 16], which have shown to be capable of solving multiple types of tasks given a particular context defined by example ‘condition-solution’ pairs. ICON is an example of a foundation model for differential equations as it is pre-trained on a variety of ODE and PDE problems, and then fine-tuned to learn specific solution operators given only example pairs of initial-boundary value data and solutions. Our probabilistic formulation of ICON provides a model for understanding the mathematics of fine-tuning of foundation models for differential equations. This ability to fine-tune ICON given new example pairs to learn solution operators of unseen differential equations is why [3, 4] describes it as an operator learner rather than an operator approximator: it adapts on-the-fly to new contexts, learning the correct operator without being given explicit information about the underlying model parameters.
@@
To understand ICON, operator learning, and foundation models of differential equations more generally, we present the notion of probabilistic operator learning, which requires modeling the data generating process that produces the training datasets. For operator learning, we model this data generating process through the formalism of random differential equations (RDEs) [17], upon which we may understand the connections of operator learning and ICON to conditional expectations, least-squares estimators, and Bayesian inference. Random differential equations provide a natural mathematical framework for understanding not just ICON, but many foundation models trained on data from many different differential equations and initial and boundary data. The probabilistic setting also provides a rigorous way of describing confidence in model predictions by characterizing model uncertainty.
@@
ICON learns a function that maps example pairs of input-output data of a differential equation along with a new unseen input, and maps it to the predicted output. This is in contrast to traditional operator learning methods in which a function maps inputs and operator parameters to solution outputs. Our primary contribution is showing that ICON directly approximates the expected value of a posterior predictive distribution, bypassing an explicit characterization of a Bayesian posterior distribution, meaning that ICON performs Bayesian inference implicitly. In other words, ICON infers the appropriate operator parameters given the example pairs and then applies the learned operator to the new inputs to predict the corresponding output in a single step. In-context learning more broadly has been understood to approximate the posterior predictive distribution [18], although this interpretation has only been studied in the context of large language models.
@@
With the knowledge that ICON approximates the mean of the posterior predictive distribution, our second main contribution is adapting ICON to generate from the posterior predictive distribution. In the context of ICON, the posterior predictive characterizes the distribution of solutions to the differential equation conditioned on example data, meaning that its statistics describes the uncertainty of the predictions as well. Tools from conditional generative modeling are needed to generate from the posterior predictive, and we show that being able to produce samples from it provides a way to quantify uncertainty in ICON’s predictions. Generating from the posterior predictive distribution is an example of conditional generative modeling. When training ICON, or foundation models for differential equations more broadly, we are given input-output pairs of conditions and solutions which come from some joint probability measure. Conditional generative modeling takes input-output pairs to characterize the distribution of outputs given unseen inputs. ICON takes this a step further where it can characterize the distribution of solutions given new conditions and input-output data that inform the model parameters.
@@ 1.1 Contributions
  • Using the RDE formalism, we explain in-context operator learning as implicitly performing Bayesian inference, in which ICON approximates the mean of the posterior predictive distribution. In Section 3, we describe in-context learning in probability spaces with Banach and Hilbert space valued random variables. Since solutions of differential equations reside in function spaces, it is natural — and indeed, necessary — to adopt this setting as model predictions will also reside in these spaces. This functional analytic framework provides the foundation for understanding when least-squares predictors coincide with conditional expectations in operator learning–based approaches to differential equations.
  • We present a generative form of in-context operator learning that is able to generate from the posterior predictive distribution. This is a form of uncertainty quantification, which describes the level of confidence ICON has in prediction. Building on the abstract formulation in Section 3, we establish the existence and properties of a generative ICON method in Section 4, as well as its ability to perform uncertainty quantification. Numerical examples that demonstrate our approach are discussed in Section 5.
@@
As we discussed in Section 1, probabilistic operator learning requires us to define the data generating process that produces the training data for operator learning. Random differential equations provides the probabilistic model that generates this training data, and provides the proper formalism for which many operator learning methods can be understood. In particular, we highlight the probability measures that are induced by random ODEs, boundary value problems, and time-dependent PDEs. Solutions to differential equations are elements of function spaces, i.e., Banach and Hilbert spaces, and therefore, the probability measures we study are defined on Banach and Hilbert spaces.
@@ 2.2 Random ODEs
where $u(t,\omega)\in\mathbb{R}$ is a function of time, $c(t,\omega)$ is a Gaussian process, and $\gamma_{1}(\omega)$, $\gamma_{2}(\omega)$, and $u_{0}(\omega)$ are $\mathbb{R}$–valued uniform random variables. The random function $a(t,u,\omega)=\gamma_{1}(\omega)c(t,\omega)+\gamma_{2}(\omega)$ is parametrized by $\gamma_{1},\gamma_{2}$, and $c(t,\omega)$. In practice, Gaussian processes are approximated by finite dimensional Gaussian vectors $\mathbf{c}\in\mathbb{R}^{\mathbf{N_{c}}}$, and so the random function $a$ is modeled by a finite dimensional random vector. The probability measure that arises from (4) can be represented in terms of a density $Q^{N}$
@@ 2.3 Random boundary value problems
Here we consider random boundary value problems (BVPs). The solution no longer evolves in time, however $V$ and $W$ are function spaces defined over some domain $E\subset\mathbb{R}^{d}$, for example, $V=L^{2}(E)$ and $W=L^{2}(\partial E)$. Generally, we may consider equations of the form
@@
where $\mathbf{u}(\omega)\in\mathbb{R}^{N}$, $\mathbf{u}_{i}(\omega)=u(x_{i},\omega)$, and $\text{d}\mathbf{W}=\,\text{d}a\,\text{d}\mathbf{c}\,\text{d}\mathbf{k}\,\text{d}u_{0}\,\text{d}u_{L}\,\text{d}\mathbf{u}$. Solutions to boundary value problems are typically approximated by finite difference or element schemes, where the solution at a particular stencil point or element is explicitly dependent only on a local neighborhood of the point or element. In this example, one would typically use a central difference scheme so that the density can be expressed with the following conditional independence structure
@@ 3.2 Probability measures of in-context operator networks
To clarify the distinction between ”classical” operator learning methods — such as DeepONets [2], Fourier neural operators [1], and neural operators more broadly [23]— and in-context operator learning, we briefly discuss the former in our probabilistic context. Classical operator learning also uses training data of the form (13) with $J=1$, but the key difference is that the operator parameters $\alpha_{m}$ are observed during training and serve as inputs to the operator. Formally, classical operator learning seeks a mapping
@@ 3.3 In-context operator learning implicitly performs Bayesian inference
ICON is an example of amortized and likelihood-free inference in a distinct way: rather than repeatedly recomputing or even explicitly characterizing the posterior distribution, it maps observations directly to the posterior predictive distribution, bypassing the bottleneck of computing the posterior entirely. This is an advantage when one simply wants a predictive model without the need to fully characterize the posterior. Posterior distributions are often difficult to characterize, particularly when they are supported on a low-dimensional manifold or are highly nonconvex, which makes sampling from them difficult. In Section 5, we consider an ODE that exhibits non-identifiability, meaning that multiple parameters may produce the same observational data. The resulting posterior then lies on a lower-dimensional manifold that is non-trivial to characterize. ICON bypasses all of these challenges by approximating the mean of the posterior predictive directly. We summarize the Bayesian explanation of deterministic in-context operator learning [3, 4] in Figure 2. In Section 4, a generative formulation of ICON will be able to characterize uncertainties of the predictive model.
@@ 4.1 Posterior predictive sampling via generative ICON
We emphasize that Theorems 4.1 and 4.2 are agnostic to the specific choice of generative modeling algorithm. While we train GenICONs using GANs in our numerical examples in Section 5, one may consider other generative models on function spaces [37, 38, 36, 35]. In Section 4.2 we provide algorithmic details describing the training of a conditional GANs from data generated for ICON.
@@
Generative ICON provides uncertainty quantification without additional cost by characterizing the distribution of predictions. By repeatedly sampling from the reference distribution with the same condition $y^{J}$ and context, GenICON produces an empirical distribution for the posterior predictive distribution from which we can estimate the predictive mean, higher moments, as well as distributions of other observables. Like ICON, GenICON is a likelihood-free and amortized inference method that can produce posterior predictive samples given context without the need to recompute the posterior distribution each time there are new observations. Figure 3 illustrates the capabilities of GenICON in comparison to standard ICON. ICON only produces the posterior predictive mean, which is a single function, while GenICON produces samples which can be used to not only characterize the mean but also the uncertainty surrounding the predictive mean. In a sense, GenICON can be viewed as a generalization of Gaussian processes, which are used to make new predictions given only a new input and data given in ordered pairs [39].
Figure 3: Illustration of the capabilities of ICON vs generative ICON. Standard ICON returns a single solution, which represents the approximation of the mean of the Bayesian posterior predictive distribution. Generative ICON characterizes uncertainty of the mean solution through samples from the posterior distribution, which is represented by the gray paths. The blue region represents the variance of the posterior predictive around the mean solution.
@@
If we are only interested in the first moment of (19), i.e. the conditional expectation (20), then Theorem 3.1 shows we do not need to consider a full probabilistic formulation, and the least squares regression (16) is sufficient. If however, we need probabilistic information, such as uncertainty quantification bounds for predictive guarantees for the conditional expectation predictions, then we need the full probabilistic formulation. The full probabilistic formulation is necessary for numerous other critical tasks, e.g. formulating and solving inverse problems using ICON, or carrying out model fine-tuning with UQ capabilities.
@@ 4.2 A generative ICON model based on generative adversarial networks
We detail one approach for constructing a generative model for sampling from the posterior predictive distribution (25). Suppose we are given the training data in the form of (16). To simplify presentation, denote the context, i.e., example condition-QoI pairs, by $\mathbf{c}=\{y^{j},z^{j}\}_{j=1}^{J-1}$. Let $\mathcal{G}_{\theta}:H\times(Y\times Z)^{J-1}\to Z$ be a family of conditional generators parametrized by $\theta$ that maps reference samples $\eta\sim\mathbb{P}_{\eta}$ and context $\mathbf{c}$ to an output sample $z^{J}_{\theta}=\mathcal{G}_{\theta}(\eta,\mathbf{c})$. The map induces a conditional measure $\mathbb{Q}_{z^{J}_{\theta}|\mathbf{c};\theta}(\cdot|\mathbf{c})=\mathcal{G}_{\theta}(\cdot,\mathbf{c})_{\sharp}\mathbb{P}_{\eta}$, as well as a parametrized family of joint measures $\mathbb{Q}_{\mathbf{c},z^{J}_{\theta};\theta}(\cdot)=\mathbb{Q}_{z^{J}_{\theta}|\mathbf{c};\theta}(\cdot|\mathbf{c})\otimes\mathbb{Q}_{\mathbf{c}}(\cdot)$, which will approximate the joint measure $\mathbb{Q}^{J}(\cdot)$ in (18).
@@ 5 Numerical examples
We demonstrate the Bayesian interpretation of ICON as well as our novel generative ICON through some simple numerical examples. Sections 5.1 and 5.2 demonstrate that ICON is robust to degenerate posteriors, model non-identifiability, and partial and noisy observations. These properties can be easily explained through a Bayesian inference. Sections 5.3 and 5.4 demonstrate generative ICON and its ability to capture the posterior predictive distribution. Examples that use ICON use the same model architecture as described in [3], while the generative ICON model architecture and training is described in Appendix A.
@@ 5.1 Robustness of ICON to degenerate posteriors and non-identifiability
We consider variations of the forward and inverse problem for Example 1 through the Bayesian interpretation of ICON. We consider the same random ODE under two different parametrizations:
@@
We let the parameters be the coefficients: $\alpha(\omega)=(\gamma_{1}(\omega),\gamma_{2}(\omega))\in\mathbb{R}^{2}$ and $\alpha(\omega)=(\eta_{1}(\omega),\eta_{2}(\omega),\eta_{3}(\omega))\in\mathbb{R}^{3}$. In the forward problem, the conditions are $y(\omega)=(c(t,\omega),u_{0}(\omega)\in L^{2}(0,T)\times\mathbb{R}$ and the quantity of interest be $z(\omega)=u(t,\omega)\in L^{2}(0,T)$. In the inverse problem, the roles are reversed — $y(\omega)=u(t,\omega)\in L^{2}(0,T)$ and $z(t,\omega)=c(t,\omega)\in L^{2}(0,T)$. For all cases, we let $T=1$. Notice that in the 2-parameter problem, the parameters are exactly identifiable with only a single demo (example-QoI pair). Given $(c(t,\omega),u(t,\omega)$, the parameters $(\gamma_{1},\gamma_{2})$ are determined by solving the linear system